OpenStudy (swaqqout_kid1):

One yard company charges $10x + $50 a month to mow and edge a lawn. A second company charges $8x + $60 a month to mow and edge a lawn. Each lawn company also offers additional services for an extra fee. Let x represent the number of additional services. Mr. Mead is comparing the two companies. How many additional services are needed for the two companies to be the same price? Mr. Mead needs additional services for the companies to cost the same price @Secret-Ninja

OpenStudy (swaqqout_kid1):

Directions: Type the correct answers in each box. Use numerals instead of words. If necessary, use / for the fraction bar(s).

OpenStudy (shinalcantara):

For them to be equal: 10x + 50 = 8x + 60 10x - 8x = 60 - 50 2x = 10 x = 5 ------------- For them to have equal amount, x=5

OpenStudy (swaqqout_kid1):

i need more if u dont mind